Marks of six students are tabulated below:

Name : Raj Anil Amit Om Rita Renu

Marks : 54 50 52 48 50 52

From the population, tabulated above, you are supposed to choose a sample of size two.

(a) Determine, how many samples of size two are possible

(b) Construct sampling distribution of means by taking samples of size 2 and organize the data.




1
Expert's answer
2021-05-12T16:11:51-0400

We have given the population,


54, 50, 52, 48, 50, 52.


Sample size "n = 2."

a.) Number of possible samples of sample size 2 without replacement "= ^6C_2 = 15"

b.)



Sampling Distribution of sample mean,

"P(X=52) = \dfrac{3}{15}"


"P(X=53) = \dfrac{2}{15}"

"P(X = 51) = \dfrac{5}{15}"

"P(X=49) = \dfrac{2}{15}"

"P(X =50) = \dfrac{3}{15}"


Mean of sampling distribution "= \dfrac{156+106+255+98+150}{15} = \dfrac{765}{15} = 51"